# Probabilistic Image-Driven Traffic Modeling via Remote Sensing

### v1

@@ 2 Related Work
Though much progress has been made, a primary challenge that remains is how to build descriptions of the environment at a global scale. Traditional approaches for modeling traffic flow assume prior knowledge of the road network and its connectivity [17], as well as the existence of large quantities of historical traffic data for each road segment [27]. As such, these methods are unable to function in locations where the road network is unknown or where road segments have insufficient historical data. To circumvent this, overhead imagery has become a useful resource for many tasks due to its dense coverage, high resolution, and ever increasing revisit rates [2, 30].
@@ 4.2 Traffic Speed Estimation
Prediction (Agg.) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Prediction (Dense) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
As each test image is associated with time-varying historical traffic data, we consider two strategies for selecting a time for evaluation. For our first experiment, we sample a time for each test image from the set of observed traffic data across all contained road segments, and use this time to generate ground-truth speed masks. We refer to this as a micro strategy, as metrics are computed globally across time. Table 1 shows the results of this study versus a recent baseline. Across all metrics, our method significantly outperforms the prior state-of-the-art.
For the second experiment, we employ a macro strategy that considers a specific set of times (Monday & Saturday, with hours 12am, 4am, 8am, 12pm, 5pm, and 8pm). For each time, we select all images in the test set that have a contained road segment with observed traffic speed data at that time. Metrics are then averaged across time such that all times are weighted equally. Table 2 shows the results of this study for a subset of times, with the overall average performance shown in the bottom row. As before, our method significantly outperforms prior work independent of the day of the week or hour of day.
@@
[45] Pseudo-Huber 10.66 8.10 0.46
Ours Pseudo-Huber 10.08 7.38 0.52
Ours Student’s t 8.94 6.70 0.62
@@
Mon (4am) 13.24 9.77 0.63 11.07 7.93 0.74
Mon (12pm) 10.41 7.87 0.52 8.85 6.61 0.65
Sat (5pm) 10.36 7.97 0.43 8.76 6.65 0.59
Sat (8pm) 10.27 7.79 0.46 8.55 6.39 0.63
Overall 11.13 8.35 0.49 9.09 6.74 0.65
@@ 4.3 Ablation Study
We conduct an extensive ablation study to validate components of our approach. First, we consider the choice of objective function. Table 1 shows the results of this study. We compare our probabilistic approach (Student’s t) versus a variant of our method which directly regresses traffic speeds using the Pseudo-Huber loss. As observed, our probabilistic approach significantly improves performance relative to this baseline. The choice of baseline objective function is consistent with that used in the prior state-of-the-art [45], highlighting that our probabilistic formulation directly leads to an increase in performance.
Next, we evaluate the impact of geo-temporal context on our traffic speed predictions. Table 3 shows the results of this experiment. All variants of our method that include geo-temporal context outperform the image-only variant. Additionally, results show that including time context is superior to including location context for predicting time-varying traffic speeds. Like-for-like comparisons with prior work show the superior performance of our approach in each setting.
We also perform an ablation study comparing strategies for integrating geo-temporal context. The results are shown in Table 4. For this experiment, we vary the location representation, context encoding, and context fusion method. The results show that using a per-pixel representation (Dense) for location is superior to the center coordinate (Center) as in prior work. Additionally, encoding location and time context using GTPE leads to better performance as compared to an MLP or CNN encoder. For fusing context features with visual features, treating location and time as a positional encoding (PE), as in GTPE, performs better than channel-wise concatenating (Concat) these features in the decoder, or adding context features as an additional token (Token) in each transformer stage. In summary, GTPE outperforms other strategies for integrating geo-temporal context, which are representative of the prior state-of-the-art.
Finally, we compare our approach, which includes the auxiliary tasks of road segmentation and orientation estimation, to variants of our approach that consider subsets of the tasks (Table 5). Results show that including the auxiliary tasks has a positive impact on performance. This matches recent work in multi-task learning that shows sharing information across tasks can lead to performance improvements when the tasks are synergistic [9].
4.4 DTS++: A Dataset for Location Adaptation
@@
[45] image 11.35 8.69 0.39
image, loc 10.95 8.44 0.43
image, time 10.68 8.12 0.46
image, loc, time 10.66 8.10 0.46
Ours image 10.34 7.74 0.49
image, loc 9.77 7.32 0.55
image, time 9.18 6.81 0.60
image, loc, time 8.94 6.70 0.62
@@
Center MLP MLP Concat Concat 10.66 8.10 0.46
Center MLP MLP Token Token 9.20 6.92 0.60
Dense CNN MLP PE Token 9.09 6.76 0.61
Dense GTPE GTPE PE PE 8.94 6.70 0.62
@@
9.25 6.91 0.60
9.11 6.75 0.61
9.01 6.74 0.62
8.94 6.70 0.62
@@
NYC Cincy original 24.11 18.89 0.03
NYC Cincy fine-tuned 12.41 9.24 0.70
@@ 5 Applications
Our approach can be thought of as estimating a local motion model from an overhead image which describes how the environment is traversed. Figure 7 visualizes how our method can be applied to novel locations to generate a local motion model. This allows our approach to be applied to novel locations where transport data is incomplete or unavailable, a common occurence. For example, road networks are not always accurate for a given location (e.g., not yet mapped or changed). Further, traffic speed data is sparse and not always available for all roads. The generalizability of our approach is made possible by the inclusion of the auxiliary tasks of road segmentation and orientation estimation, in addition to their performance benefit from multi-task learning.
Figure 6: Visualizing traffic speed predictions for individual road segments vs time. (left) A road segment shown in red. (right) Our approach’s predictions (red) versus historical traffic data (green).
@@ 6 Conclusion
Probabilistic Image-Driven Traffic Modeling via Remote Sensing
@@
1 Extended Evaluation: iNaturalist 2018
We investigate the performance of our geo-temporal positional encoding (GTPE) on a different task, fine-grained visual categorization, using the iNaturalist 2018 dataset. This dataset contains over $450,000$ images of birds (approx. $8000$ unique species) collected from the iNaturalist social network. The task is to perform large-scale, fine-grained species classification in the scenario where there are many visually similar species and high class imbalance.
For this experiment, we take advantage of the framework introduced by Aodha et al. [31]. The authors proposed to learn a spatiotemporal prior, $P(y|\boldsymbol{x})$, conditioned on location and time context, $\boldsymbol{x}$, that estimates the probability that a given object category, $y$, occurs at that location. The spatiotemporal prior is combined with the output of an image classifier, $P(y|I)$. The results show that including the spatiotemporal prior helps disambiguate fine-grained categories. We take advantage of this framework, but use our proposed geo-temporal positional encoding (GTPE) module to learn the spatiotemporal prior, instead of their approach which uses a MLP consisting of several residual layers and a final output embedding layer (each of 256 dimensions).
@@ 1 Extended Evaluation: iNaturalist 2018
2 Visualizing Geo-Temporal Context
@@
January $May$
September
Refer to caption Refer to caption Refer to caption
Monarch Butterfly
Refer to caption Refer to caption Refer to caption
Western Honey Bee
Our method integrates location and time through distinct pathways in GTPE which can be thought of as 64-dimensional embeddings. To visualize the learned feature representations for time, we construct a false color image using principal component analysis (PCA), shown in Figure S1. For every time combination (day of week, time of day), we extract its feature representation from the time embedding, and stack them. This results in a matrix of size $168\times 64$. We apply PCA to reduce the feature dimensionality and reshape to $7\times 24$ (days by hours). The resulting image represents time-dependent traffic patterns learned by our approach.
@@
3 Dynamic Traffic Speeds++ (DTS++)
We introduced DTS++, a new dataset for image-driven traffic modeling that includes a new city, Cincinnati, OH. Figure S3 gives a visual overview of the coverage of the dataset (traffic speeds averaged across time). It includes $11,137$ overhead images ($1024\times 1024$) at approximately $0.3$ meters per-pixel, paired with one year of historical traffic data (2018) collected from Uber Movement Speeds [1]. In Figure S4, we visualize the average road segment speed versus time for individual days of the week. Additionally, Figure S5 shows the average road segment speed for different types of roads using OpenStreetMap’s road type classification. Compared to New York City (from DTS), Cincinnati has very different spatiotemporal mobility patterns. For example, the maximum observed traffic speed in Cincinnati for a motorway is on the order of 100 km/h, where in New York City it is 80 km/h.
Figure S3: Example traffic speed data from DTS++ for Cincinnati (visualized as average speeds).
Figure S4: Visualizing the average road segment speed over time in Cincinnati for different days.
4 Capturing Uncertainty in Traffic Speeds
@@
Figure S5: Visualizing the average road segment speed against time in Cincinnati, where each road is categorized by its OpenStreetMap road type classification.
@@
Image Road (Label) Road (Pred.) Orientation (Label) Orientation (Pred.)
@@
Figure S7: Qualitative results for the auxiliary tasks of road segmentation and orientation estimation. Orientation ($-\pi$ to $\pi$) is represented by $K=16$ bins using the HSV color map.
@@ 7 Detailed Architecture
Table S2: Evaluation of auxiliary tasks.
Speed Road Orientation
(RMSE $\downarrow$) (F1 Score $\uparrow$) (Accuracy $\uparrow$)
8.94 77.51% 73.78%

### v2

@@ Abstract
Keywords:
Traffic Modeling Geo-Temporal Context Remote Sensing
@@ 2 Related Work
Though much progress has been made, a primary challenge that remains is how to build descriptions of the environment at a global scale. For example, traditional approaches for modeling traffic flow assume prior knowledge of the road network and its connectivity [18], as well as the existence of large quantities of historical traffic data for each road segment [28]. As such, these methods are unable to function in locations where the road network is unknown or where road segments have insufficient historical data. To circumvent issues of scale, overhead imagery has become a useful resource for many tasks due to its dense coverage, high resolution, and ever increasing revisit rates [2, 31].
@@ 4.2 Traffic Speed Estimation
As each test image is associated with time-varying historical traffic data, we consider two strategies for selecting a time for evaluation. For our first experiment, we sample a time for each test image from the set of observed traffic data across all contained road segments, and use this time to generate a ground-truth speed mask. We refer to this as a micro strategy, as metrics are computed globally across time. Table 2 shows the results of this study. Across all metrics, our method significantly outperforms the prior state-of-the-art.
For the second experiment, we employ a macro strategy that considers a specific set of times (Monday & Saturday, with hours 12am, 4am, 8am, 12pm, 5pm, and 8pm). For each time, we select all images in the test set that contain a road segment with observed traffic speed data at that time. Metrics are then averaged across time such that all times are weighted equally. Table 2 shows the results of this study for a subset of times, with the overall average performance shown in the bottom row. As before, our method significantly outperforms prior work independent of the day of the week or hour of day.
@@
[47] Pseudo-Huber $10.64$ $8.09$ $0.46$
Ours Pseudo-Huber $10.02$ $7.34$ $0.52$
Ours Student’s t $8.84$ $6.64$ $0.63$
@@
Mon (4am) $13.20$ $9.74$ $0.63$ $10.95$ $7.85$ $0.75$
Mon (12pm) $10.40$ $7.86$ $0.52$ $8.78$ $6.56$ $0.66$
Sat (5pm) $10.35$ $7.96$ $0.43$ $8.68$ $6.60$ $0.60$
Sat (8pm) $10.26$ $7.78$ $0.46$ $8.47$ $6.34$ $0.63$
Overall $11.12$ $8.34$ $0.49$ $9.15$ $6.76$ $0.65$
Pred. (Agg.) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Pred. (Dense) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 4.3 Ablation Study
We conduct an extensive ablation study to validate components of our approach. First, we consider the choice of objective function. Table 2 shows the results of this experiment. We compare our probabilistic approach (Student’s t) versus a variant of our method which directly regresses traffic speeds using the Pseudo-Huber loss. As observed, our probabilistic approach significantly improves performance relative to this baseline. The choice of baseline objective function is consistent with that used in the prior state-of-the-art [47], highlighting that our probabilistic formulation directly leads to an increase in performance.
Next, we evaluate the impact of geo-temporal context on our traffic speed predictions. Table 3 shows the results of this experiment. First, an image-only variant of our approach outperforms a metadata-only variant (loc, time), demonstrating the importance of visual features for this task. Further, adding both location and time context leads to additional performance gains, with time context being superior to location context for predicting time-varying traffic speeds. Finally, like-for-like comparisons with prior work show the superior performance of our approach in each setting.
@@
[47] loc, time $13.13$ $9.60$ $0.18$
image $11.33$ $8.68$ $0.39$
image, loc $10.94$ $8.43$ $0.43$
image, time $10.67$ $8.11$ $0.46$
image, loc, time $10.64$ $8.09$ $0.46$
Ours loc, time $12.16$ $9.20$ $0.30$
image $10.28$ $7.70$ $0.50$
image, loc $9.65$ $7.26$ $0.56$
image, time $9.10$ $6.75$ $0.61$
image, loc, time $8.84$ $6.64$ $0.63$
@@
Center MLP MLP Concat Concat $10.64$ $8.09$ $0.46$
Center MLP MLP Token Token $9.13$ $6.87$ $0.61$
Dense CNN MLP PE Token $9.02$ $6.71$ $0.61$
Dense GTPE GTPE PE PE $8.84$ $6.64$ $0.63$
We also perform an ablation study comparing strategies for integrating geo-temporal context. The results are shown in Table 4. For this experiment, we vary the location representation, context encoding, and context fusion method. The results show that using a per-pixel representation (Dense) for location is superior to the center coordinate (Center) as in prior work [47]. Additionally, encoding location and time context using GTPE leads to better performance as compared to an MLP or CNN encoder. For fusing context features with visual features, treating location and time as a positional encoding (PE), as in GTPE, performs better than channel-wise concatenating (Concat) these features in the decoder, or adding context features as an additional token (Token) in each transformer stage (e.g., [12]). In summary, GTPE outperforms other strategies for integrating geo-temporal context, which are representative of the prior state-of-the-art.
@@
$9.18$ $6.86$ $0.60$
$9.03$ $6.70$ $0.61$
$8.92$ $6.67$ $0.62$
$8.84$ $6.64$ $0.63$
Finally, we compare our approach, which includes the auxiliary tasks of road segmentation and orientation estimation, to variants of our approach that consider subsets of the tasks (Table 5). Results show that including the auxiliary tasks has a positive impact on performance. This matches recent work in multi-task learning that finds sharing information across tasks can lead to performance improvements when the tasks are synergistic [9].
4.4 DTS++: A Dataset for Location Adaptation
@@
NYC Cincy original $24.06$ $18.84$ $0.04$
NYC Cincy fine-tuned $12.29$ $9.18$ $0.75$
@@ 5 Applications
Figure 6: Visualizing traffic speed predictions for individual road segments versus time. (left) A road segment shown in red. (right) Predictions from our approach (red) versus historical traffic data (green).
@@
Our approach can be thought of as estimating a local motion model from an overhead image which describes how the environment is traversed. Figure 7 visualizes how our method can be applied to novel locations to generate a local motion model. This allows our approach to be applied to locations where transport data is incomplete or unavailable, a common occurrence. For example, road networks are not always accurate for a given location (e.g., unmapped or changed). Further, traffic speed data is sparse and not always available for all roads. The generalizability of our approach is made possible by the inclusion of the auxiliary tasks of road segmentation and orientation estimation, in addition to their performance benefit from multi-task learning.
@@ 6 Conclusion
Probabilistic Image-Driven Traffic
Modeling via Remote Sensing
@@
1 Dynamic Traffic Speeds++ (DTS++)
Figure S1: Example traffic speed data from DTS++ for Cincinnati (road segment speeds averaged over time). Green (red) corresponds to fast (slow) traffic speeds.
We introduce DTS++, an extension of the Dynamic Traffic Speeds (DTS) dataset [47] to include an additional city, Cincinnati, OH. To construct DTS++, we emulated the process used to create DTS, which we outline here. First, we obtained a year of historical traffic speed data (2018) from Uber Movement Speeds [1], a dataset of publicly available aggregated speed data derived from Uber rideshare trips. The speed data is provided as averages over road segments at an hourly resolution, with the underlying road network described using OpenStreetMap. Figure S1 gives a visual overview of Cincinnati’s spatial coverage (traffic speeds averaged across time).
To align the historical traffic speed data with overhead imagery, we first generated a set of non-overlapping spherical Mercator tiles (XYZ style) that intersect a bounding box around Cincinnati. We identified tiles containing road segments with available speed data, and for each, downloaded an overhead image from Bing Maps (at a resolution of approximately $0.3$ meters per-pixel). This process resulted in $11,137$ overhead images ($1024\times 1024$) with associated road geometries and traffic speed data. To support model training and evaluation, these non-overlapping tiles are randomly split into 85% training, 5% validation, and 10% testing, such that road segments observed during testing are not observed during training.
In Figure S2 (left), we visualize the average road segment speed versus time for individual days of the week. Additionally, Figure S2 (right) shows the average road segment speed for different types of roads using OpenStreetMap’s road type classification. DTS++ extends the DTS dataset by adding a new city, Cincinnati, essentially doubling the size of the dataset and providing support for location adaptation experiments. Compared to New York City (from DTS), Cincinnati has very different appearance characteristics as well as spatiotemporal mobility patterns. For example, the maximum observed traffic speed in Cincinnati for a motorway is $\sim 100$ km/h, where in New York City it is $\sim 80$ km/h.
Figure S2: (left) Visualizing the average road segment speed over time in Cincinnati for different days. (right) Visualizing the average road segment speed against time in Cincinnati, where each road is categorized by its OpenStreetMap road type classification.
2 Fine-Grained Visual Categorization
{tblr}
colspec=X[1,c,m] *3X[2,c,m],width=1rowsep=1pt,colsep=5pt,vspan=even,
& January May September
[…]
Monarch Butterfly
[…]
Western Honey Bee
We investigate the performance of our geo-temporal positional encoding (GTPE) module on a different task, fine-grained visual categorization, using the iNaturalist 2018 dataset [43]. This dataset contains over $450,000$ images of birds (approximately $8000$ unique species) collected from the iNaturalist social network. The task is to perform large-scale, fine-grained species classification in the scenario where there are many visually similar species and high class imbalance.
For this experiment, we take advantage of the framework introduced by Aodha et al. [32]. The authors proposed to learn a spatiotemporal prior, $P(y|\boldsymbol{x})$, conditioned on location and time context, $\boldsymbol{x}$, that estimates the probability that a given object category, $y$, occurs at that location. The spatiotemporal prior is combined with the output of an image classifier, $P(y|I)$. The results show that including the spatiotemporal prior helps disambiguate fine-grained categories. We take advantage of this framework, but use our proposed geo-temporal positional encoding module to learn the spatiotemporal prior, instead of their approach which uses a MLP consisting of several residual layers and a final output embedding layer (each of 256 dimensions).
@@ 2 Fine-Grained Visual Categorization
3 Additional Experiments
3.1 Extended Ablation
To evaluate our choice of backbone architecture, we trained a variant of our approach replacing our encoder (which combines convolutional layers and attention layers) with the hierarchical transformer encoder from SegFormer [51], omitting contextual inputs (location and time). This is equivalent to our image-only variant. When evaluated on New York City, the performance of this baseline is 10.51 RMSE, 7.79 MAE, and 0.48 $R^{2}$, which is worse than our image-only variant (10.28, 7.70, and 0.50 respectively), but inline with expectations. Though these results show our choice of backbone architecture to be superior for image-driven traffic modeling, we would note that our proposed components (e.g., GTPE, probabilistic formulation) could conceivably be combined with any backbone architecture.
3.2 Visualizing Geo-Temporal Context
@@
Our method integrates location and time through distinct pathways in the GTPE module which can be thought of as 64-dimensional embeddings. To visualize the learned feature representations for time, we construct a false color image using principal component analysis (PCA), shown in Figure S4. For every time combination (day of week, hour of day), we extract its feature representation from the time embedding, and stack them. This results in a matrix of size $168\times 64$. We apply PCA to reduce the feature dimensionality and reshape to $7\times 24$ (days by hours). The resulting image represents time-dependent traffic patterns learned by our approach.
@@
3.3 Capturing Uncertainty in Traffic Speeds
@@
Image Road (Label) Road (Pred.) Orient. (Label) Orient. (Pred.)
@@
Figure S6: Qualitative results for the auxiliary tasks of road segmentation and orientation estimation. Orientation is represented by $K=16$ bins using the HSV color map.
@@ 3.4 Auxiliary Tasks
Table S2: Evaluation of auxiliary tasks.
Speed Road Orientation
(RMSE $\downarrow$) (F1 Score $\uparrow$) (Accuracy $\uparrow$)
8.84 77.51% 73.78%
3.5 Complexity Analysis
Our full method has approximately 18.1 million trainable parameters and the total estimated model parameters size is 72.57 MB. It takes about 48 hours to train on a single NVIDIA A6000 (for fair comparison with baselines, we use a batch size of two and accumulate gradients across eight batches). We estimate that our approach has 1587.21 GFLOPS (floating point operations) and 792.41 GMACs (multiply-add operations). During inference, our unoptimized implementation achieves approximately 5.2 frames per second (batch size of 1).